A box has two coins. One has two heads and the other is a fair coin with one head and one tail. Now you flip both coins at the same time.

a.) How many sample points are there in the sample space? List all the sample points.

b.) What is the probability that you come up with two heads?
Answer #1

a) For first coin, we have only one choice and for second coin, we have two choices. So,

Number of sample points = 1*2 = 2

Sample points will be:

{HH, HT}

b) P(Two heads)

= P(HH)

= 1/2